OpenStudy (anonymous):

is this correct? solve for x 2(y-3)=-8(x-2) (-y=22)/4=x

OpenStudy (ghazi):

2(y-3)=-8(x-2) 2y-6=-8x+16 2y=-8x+22 divide by two on both sides y=-4x+11

OpenStudy (anonymous):

ii need to solve for x

OpenStudy (anonymous):

\[2(y-3)=-8(x-2) \rightarrow -\frac{1}{4}(y-3)=x+2 \rightarrow x=-\frac{1}{4}(y-3)-2\]

OpenStudy (ghazi):

x=-(y-11)/4

OpenStudy (anonymous):

Sorry, typo. should be +2 at the end there.

OpenStudy (anonymous):

\[2(y−3)=−8(x−2)→−\frac{1}{4}(y−3)=x-2→x=−\frac{1}{4}(y−3)+2.\] Then you can simplify from there.

OpenStudy (precal):

from here you could solve for x